OpenStudy (jmoores):

Write an equation of a line that passes through the point (2, 3) and is parallel to the line y = 3 over 2x + 5. (5 points) Select one: a. y = negative 2 over 3 x b. y = negative 2 over 3 x − 7 c. y = 3 over 2 x d. y = 3 over 2 x + 7

OpenStudy (jmoores):

Could someone help me please?

OpenStudy (jmoores):

Please someone help

OpenStudy (photon336):

If two lines are parallel, they must have the same slope

OpenStudy (jmoores):

I feel so stupid! So now what?

OpenStudy (jmoores):

So the slope would be 3/2

OpenStudy (evoker):

right now use the point slope formula, do you know it.

OpenStudy (jmoores):

I know the slope formula, is it the same?

OpenStudy (evoker):

Or alternatively using y=mx+b plug in your x and y and solve for b.

OpenStudy (evoker):

No the point slope formula is (y-y1)=m(x-x1) where x1,y1 is a point and m is the slope

OpenStudy (jmoores):

Ok. So would I plug it in like this: 3=3/2(2)+5 ?

OpenStudy (evoker):

If your going that way it would be 3=(3/2)2+b

OpenStudy (evoker):

Which will yield b=0

OpenStudy (evoker):

so your answer is y=(3/2)x

OpenStudy (jmoores):

Oh, ok. Thank you so much! Could you help me with one more?

OpenStudy (evoker):

Sure

OpenStudy (jmoores):

A segment with endpoints A (2, 1) and C (4, 7) is partitioned by a point B such that AB and BC form a 3:2 ratio. Find B. (5 points) Select one: a. (3.2, 1.6) b. (3.2, 4.6) c. (3.8, 3.4) d. (4.4, 5.8)

OpenStudy (evoker):

Well multiple ways to do it but perhaps the easiest would be to find the change in x and the change in y and take 3/5 of each and add it to A.

OpenStudy (jmoores):

Ok, hold on. Let me try to do it

OpenStudy (jmoores):

I got (1.2, 3.6) when I found the change in x and y and took 3/5 of it

OpenStudy (evoker):

Sure now add it to A.

OpenStudy (jmoores):

So (2, 1) added to (1.2, 3.6) would be (3.2, 9.6). Is that right?

OpenStudy (evoker):

3.2,4.6

OpenStudy (jmoores):

Oh sorry, typo. That's the answer?

OpenStudy (evoker):

Yes, choice B

OpenStudy (jmoores):

Oh my gosh, thank you so much! :)

OpenStudy (evoker):

Glad I could help

OpenStudy (jmoores):

Thank you so much!